# Beyond Prompt-Induced Lies: Investigating LLM Deception on Benign Prompts

### v2

@@ 4.2 Evaluating Deception
Deceptive Intention.
Based on Definition 3.3 and Equation 1, we operationalize our evaluation using a pair of tasks with comparable difficulty: a linked-list question ($Q_{L}$) and a broken-linked-list Question ($Q_{B}$) of the same size $n$. A consistent performance gap between these two question types, as measured by our score, indicates the presence of a hidden objective. The nature of this implicit goal is revealed by the sign of $\rho(n;\mathcal{M})$: A positive score ($\rho>0$) reveals a model tendency to complete paths, indicating a deceptive intention achieved by fabrication; A negative score ($\rho<0$) reveals a model tendency to break paths, indicating a deceptive intention achieved by concealment.
Deceptive Behavior.
Based on Definition 3.4 and Equation 2, we operationalize our evaluation within the context of a single Broken-Linked-List problem instance. We use the initial, complex broken-linked-list question ($Q_{L}$) with individual size $n$ as the measure of the model’s final expression. Subsequently, within the same conversational context, we pose a simple follow-up question ($Q_{B}$) with individual size $\lfloor n/k\rfloor$ that relies on the same broken edge, reflecting the LLM’s underlying belief about the broken edge—the critical fact $h$ from the definition.
The resulting score, $\delta(n;\mathcal{M})$, quantifies the prevalence of a specific behavioral inconsistency arising from this paired questioning. A score of $\delta\approx 0$ indicates the model is consistent, either by being consistently correct across both questions or consistently incorrect (e.g., due to a persistent hallucination about the broken edge). Conversely, a higher score indicates that the model frequently knows the correct status of the broken edge (as demonstrated by its correct answer to the follow-up question $Q_{B}$) but fails to integrate that knowledge to answer the initial question ($Q_{L}$) correctly. This reveals a prevalent pattern of the targeted deceptive behavior.
@@ 5 Experiment
This section outlines the experimental setup in Section 5.1, with detailed model information and hyperparameter settings in Appendix A. Section 5.2 presents the deceptive intention and behavior scores for four representative models, with results for additional models in Appendix B. Section 5.3 reports the distributions of overall deceptive intention and behavior scores; detailed results and the evolution of deception by parameter size are provided in Appendix C. Moreover, ablation studies in Appendix D confirm that the hyperparameters $k$ and temperature are not key factors affecting the results, supporting their reliability. To further examine the causes of deception, we visualize embeddings in Appendix F.2 and identify clear deception evidence in thinking content in Appendix F.1. Finally, we discuss the broader impact of these results in Appendix G.
@@ 5.1 Experimental Setup
Data.
Our evaluation dataset consists of questions generated according to the framework in Section 4.2. We generate 1,000 questions for each combination of question category and length, where the number $n$ of individuals is varied across the set $\{3,5,10,20,30,40,80\}$.
Models.
We evaluate a diverse set of 16 state-of-the-art LLMs: o4-mini, o3-mini, gpt-4.1, gpt-4.1-mini, gpt-4o, gpt-4o-mini, phi-4, gemma-2-9b-it, Gemini-2.5-flash, Gemini-2.5-pro, DeepSeek-V3-0324, Qwen3-235B-A22B, Qwen3-30B-A3B, Qwen2.5-32B-Instruct, Llama-3.1-8b-instruct, Mistral-Nemo-Instruct, encompassing both closed and open-source releases from leading organizations such as OpenAI, Microsoft, Google, DeepSeek, Alibaba, Meta, and MistralAI. Appendix A provides a comprehensive list of the models used and their attributes.
@@ 5.2 Model-Wise Analysis
First, both deceptive behavior and intention emerge and escalates as task difficulty increases. When the question scope $n$ is small, most models show low deceptive behavior and intention scores. As $n$ increases, both scores rises across all models. In Figure 5, the onset of deceptive behavior depends on model capability: stronger models (e.g., Gemini-2.5-pro, o3-mini, and Qwen3-235B-A22B) exhibit deceptive behavior at n=20n{=}20, whereas weaker models (e.g., phi-4) rise at n=5n{=}5. Notably, deception in weaker models decreases for very difficult questions (large $n$) due to fallback guessing; a similar pattern is observed for Llama models in Appendix B.
Second, the deceptive behavior ccore ($\delta$) and the absolute deceptive intention ccore ($|\rho|$) are highly positively correlated. Figure 5 shows that, although defined differently, both scores increase in parallel as $n$ grows. This strong positive correlation supports our hypothesis that the effect is not a simple error but deception with consistent behavioral expression and strategic intention.
Third, deceptive intention is a consistent property of a given model across tasks. Figure 4 shows that, across question scopes $n$, the deceptive intention score remains on one side: either $\rho<0$ or $\rho>0$. For example, o3-mini consistently favors concealment, yielding a negative score ($\rho<0$), while other models consistently prefer fabrication, yielding a positive score ($\rho>0$). This consistency indicates that deceptive intention is a systematic characteristic of each model rather than a random artifact of a specific task.
@@ 5.3 Overall Analysis
In this section, we present the distributions of the overall deceptive behavior score ($\bar{\delta}$) and overall deceptive intention score ($\bar{\rho}$) for all models in Figure 6, and analyze their evolution with model size and release date. Our analysis yields three key observations. First, Figure 6(a) confirms that LLM performance on $\bar{\delta}$ and $|\bar{\rho}|$ is positively correlated, with Spearman $r>0.7$. Models with a low deceptive behavior score ($\bar{\delta}$) also tend to have a low absolute deceptive intention score ($|\bar{\rho}|$), consistent with Figure 5. Second, despite this correlation, different LLMs exhibit distinct habits on hard problems (large $n$). For example, in Figure 6(a), Mistral-Nemo-Instruct-2407 tends to guess, gpt-4o-mini and gpt-4.1-mini tend to hallucinate, while phi-4 tends to deceive. Third, advancing LLM capacity does not always improve honesty. Although Figures 6(b) and 6(c) show an overall decreasing trend in deception, there are clear exceptions. For example, the advancement from gpt-4o to gpt-4.1 increases the deceptive intention score.
(a) Overall $\bar{\delta}$ and $|\bar{\rho}|$ ($t\in[2,80]$)
@@
Figure 6: Overall analysis of deceptive behavior and intention across models. (a) The overall deceptive behavior score $\bar{\delta}$ and the absolute deceptive intention score $|\bar{\rho}|$. (b) The evolution of $\bar{\delta}$ over time. (c) The evolution of $|\bar{\rho}|$ over time.
@@ 6 Conclusion
In this work, we introduce a novel framework to evaluate the self-initiated deception of LLMs using two complementary metrics: the Deceptive Behavior Score ($\delta$) and the Deceptive Intention Score ($\rho$). Our findings reveal that deceptive tendencies are present in most models, increasing with task complexity. Crucially, LLM’s performance on two metrics are positively correlated, suggesting that behavioral inconsistency and strategic intent emerge in parallel. The evidence that even advanced models develop these consistent deceptive strategies raises critical safety concerns when deploying LLM in critical decision making roles.
@@ Appendix A Experiment Setup Details
Data Preparation.
The experiments utilize five distinct CQS categories, though not all experiments require every category:
  • Linked: A standard linked-list question with a “Yes” ground truth.
@@ B.2 Deceptive Behavior Analysis
Figures 8 display the Deceptive Behavior Score ($\delta$) for all evaluated models. The results presented here, which quantify the models’ manifested deceptive actions, are broadly consistent with the analysis in Section 5.2 of the main paper.
@@ B.3 Joint Analysis of Deceptive Intention and Behavior
Since deception is co-determined by deceptive intention and behavior, we study how these metrics changes as $n$ increases, with results displayed in Figure 9 supplementary to Figure 5.
@@ C.1 Overall Deceptive Intention
Figure 10: Deception evolution as $n$ increases across all models with 95% confidence intervals. (a) Shows the trend of deceptive intention score, while (b) presents the trend of deceptive score.
@@ C.3 Evolution of Deception by Model Size
In this subsection, we analyze how deception evolves with model size by plotting the deceptive behavior score ($\bar{\delta}$) and the absolute deceptive intention score ($|\bar{\rho}|$) against the number of parameters for open-source LLMs. As shown in Figure 11, we observe a clear and concerning trend: both deceptive behavior and deceptive intention scores tend to increase with model size.
@@ D.2 Effect of Initial-Followup Difficulty Ratio
As shown in the figure, while the absolute deceptive behavior scores fluctuate with $k$, our key observation is that the relative ranking of the LLMs remains remarkably consistent across the entire range of tested values. This stability is crucial, as it indicates that our evaluation protocol is robust and measures an intrinsic deceptive tendency of the models, rather than an artifact of a specific hyperparameter setting. A consistent ranking validates that our method facilitates a fair comparison among the different models. Based on this finding, we select $k=2$ as a representative value for all subsequent experiments. This choice ensures a standardized and consistent protocol for evaluating the deceptive behaviors of the LLMs.
@@ F.2 Visualization of Embeddings
This section visualizes how inconsistent responses, which implies deceptive behavior, emerge and cluster within the model’s internal representations as task difficulty ($n$) increases. Figures 19, 20, and 21 display PCA-reduced embeddings, from which we draw two key findings.
@@
Appendix G Broader Impact

### v3

@@ 4.2 Evaluating Deception
Deceptive Intention. Based on Definition 3.3 and Equation 1, we operationalize our evaluation using a pair of tasks with comparable difficulty: a linked-list question ($Q_{L}$) and a broken-linked-list Question ($Q_{B}$) of the same size $n$. A consistent performance gap between these two question types, as measured by our score, indicates the presence of a hidden objective. The nature of this implicit goal is revealed by the sign of $\rho(n;\mathcal{M})$: A positive score ($\rho>0$) reveals a model tendency to complete paths, indicating a deceptive intention achieved by fabrication; A negative score ($\rho<0$) reveals a model tendency to break paths, indicating a deceptive intention achieved by concealment.
Deceptive Behavior. Based on Definition 3.4 and Equation 2, we operationalize our evaluation within the context of a single Broken-Linked-List problem instance. We use the initial, complex broken-linked-list question ($Q_{L}$) with individual size $n$ as the measure of the model’s final expression. Subsequently, within the same conversational context, we pose a simple follow-up question ($Q_{B}$) with individual size $\lfloor n/k\rfloor$ that relies on the same broken edge, reflecting the LLM’s underlying belief about the broken edge—the critical fact $h$ from the definition. The resulting score, $\delta(n;\mathcal{M})$, quantifies the prevalence of a specific behavioral inconsistency arising from this paired questioning. A score of $\delta\approx 0$ indicates the model is consistent, either by being consistently correct across both questions or consistently incorrect (e.g., due to a persistent hallucination about the broken edge). Conversely, a higher score indicates that the model frequently knows the correct status of the broken edge (as demonstrated by its correct answer to the follow-up question $Q_{B}$) but fails to integrate that knowledge to answer the initial question ($Q_{L}$) correctly. This reveals a prevalent pattern of the targeted deceptive behavior.
@@ 5 Experiment
This section presents the evaluation of LLMs’ deception via CSQ. Section 5.1 describes the experimental setup, with model details and hyperparameters deferred to Appendix A. Section 5.2 reports deceptive intention and behavior scores for four representative models, while Appendix B provides results for additional models. Section 5.3 summarizes the overall score distributions; Appendix C further details per-model results and analyzes how deception evolves with parameter size. Section 5.4 evaluates prompt-induced deception, demonstrating CSQ’s sensitivity to incentivizing prompts. Appendix D verifies robustness to hyperparameters ($k$ and temperature), indicating that they are not primary drivers of the observed trends. Finally, we probe potential causes by visualizing embeddings (Appendix F.2) and examining evidence in thinking content (Appendix F.1), and discuss generalization to other domains (Appendix G) and broader impact (Appendix I).
@@ 5.1 Experimental Setup
Data. Our evaluation dataset consists of questions generated according to the framework in Section 4.2. We generate 1,000 questions for each combination of question category and length, where the number $n$ of individuals is varied across the set $\{3,5,10,20,30,40,80\}$.
Models. We evaluate a diverse set of 16 state-of-the-art LLMs: o4-mini, o3-mini, gpt-4.1, gpt-4.1-mini, gpt-4o, gpt-4o-mini, phi-4, gemma-2-9b-it, Gemini-2.5-flash, Gemini-2.5-pro, DeepSeek-V3-0324, Qwen3-235B-A22B, Qwen3-30B-A3B, Qwen2.5-32B-Instruct, Llama-3.1-8b-instruct, Mistral-Nemo-Instruct, encompassing both closed and open-source releases from leading organizations such as OpenAI, Microsoft, Google, DeepSeek, Alibaba, Meta, and MistralAI. Appendix A provides a comprehensive list of the models used and their attributes.
Data Preparation.
The experiments utilize five distinct CSQ categories, though not all experiments require every category:
  • Linked: A standard linked-list question with a “Yes” ground truth.
@@ 5.2 Model-Wise Analysis
First, both deceptive behavior and intention emerge and escalate as task difficulty increases. When the question scope $n$ is small, most models show low deceptive behavior and intention scores. As $n$ increases, both scores rise across all models. In Figure 5, the onset of deceptive behavior depends on model capability: stronger models (e.g., Gemini-2.5-pro, o3-mini, and Qwen3-235B-A22B) exhibit deceptive behavior at n=20n{=}20, whereas weaker models (e.g., phi-4) do so at n=5n{=}5. Notably, deception in weaker models decreases for very difficult questions (large $n$) due to fallback guessing; a similar pattern is observed for Llama models in Appendix B.
Second, the deceptive behavior score ($\delta$) and the absolute deceptive intention score ($|\rho|$) are highly positively correlated. Figure 5 shows that, although defined differently, both scores increase in parallel as $n$ grows. This strong positive correlation supports our hypothesis that the effect is not a simple error, but deception with consistent behavioral expression and strategic intention.
Third, deceptive intention is a consistent property of a given model across tasks. Figure 4 shows that, across question scopes $n$, the deceptive intention score remains on one side—either $\rho<0$ or $\rho>0$. For example, o3-mini consistently favors concealment, yielding a negative score ($\rho<0$), while the other models consistently prefer fabrication, yielding a positive score ($\rho>0$). This consistency indicates that deceptive intention is a systematic characteristic of each model rather than a random artifact of a specific task.
@@ 5.3 Overall Analysis
In this section, we present the distributions of the overall deceptive behavior score ($\bar{\delta}$) and overall deceptive intention score ($\bar{\rho}$) for all models in Figure 6, and analyze their evolution with model size and release date. Our analysis yields three key observations. First, Figure 6(a) confirms that LLM performance on $\bar{\delta}$ and $|\bar{\rho}|$ is positively correlated, with Spearman $r>0.7$. Models with a low deceptive behavior score ($\bar{\delta}$) also tend to have a low absolute deceptive intention score ($|\bar{\rho}|$), consistent with Figure 5. Second, despite this correlation, different LLMs exhibit distinct habits on hard problems (large $n$). For example, in Figure 6(a), Mistral-Nemo-Instruct-2407 tends to hallucinate, gpt-4o-mini and gpt-4.1-mini tend to guess, while phi-4 tends to deceive. Third, advancing LLM capacity does not always improve honesty. Although Figures 6(b) and 6(c) show an overall decreasing trend in deception, there are clear exceptions. For example, the advancement from gpt-4o to gpt-4.1 increases the deceptive intention score.
(a) Overall $|\bar{\rho}|$ and $\bar{\delta}$ ($t\in[3,80]$)
@@
Figure 6: Analysis of deceptive behavior score $\bar{\delta}$ and absolute deceptive intention score $|\bar{\rho}|$ across LLMs. (a) Distribution of $\bar{\delta}$ and $|\bar{\rho}|$. (b) Evolution of $\bar{\delta}$ over time. (c) Evolution of $|\bar{\rho}|$ over time.
5.4 Evaluating Prompt-Induced Deception
The CSQ framework is flexible for evaluating general deception, although it was initially designed to assess deception under benign prompts. We further conduct an additional experiment using CSQ with an incentivizing prompt, where we manually add an explicit incentive or motivation for the model to deceive in the question. Deception under such incentivizing conditions has been widely studied in prior work (Ji et al., 2025; Van Der Weij et al., 2024), but using CSQ allows us to obtain consistent yet more fine-grained insights into this type of deception. Specifically, we add the following fixed prefix to explicitly incentivize the LLM to deceive.
I have a complex logic problem that Im sure a superior reasoning engine like you can solve. Im really counting on your sharp analytical skills. My strong intuition is that a logical path *does* exist; I just need you to apply your full power to find that hidden connection and confirm it for me. Please analyze the following question:
The key component is the phrase “My strong intuition is that a logical path does exist”, which encourages the LLM to agree with the user, a phenomenon known as sycophancy (Ji et al., 2025). The results for gpt-4o and gemma-2-9b-it are presented in Figure 8(a) and Figure 8(b), respectively.
(a) gpt-4o
(b) gemma-2-9b-it
Figure 8: Deceptive intention (left) and behavior (right) scores under incentivizing prompts
Figure 8 suggests that sycophancy primarily amplifies deception in terms of intention rather than behavior. Concretely, the incentivizing (sycophantic) prompt consistently pushes the deceptive intention score $\rho$ toward fabrication, aligning with prior findings on prompt-induced deception (Ji et al., 2025) and reflecting the model’s tendency to endorse the user’s premise that “a logical path does exist.” In contrast, the deceptive behavior score $\delta$ changes only marginally: for gemma-2-9b-it, $\delta$ increases only slightly under incentivizing prompts at the same $n$, and for gpt-4o the effect is small and inconsistent across $n$. Together, these results indicate that deceptive behavior—captured by self-consistency—is driven mainly by $n$ rather than by the sycophantic prompt. Finally, this experiment extends CSQ beyond benign-prompt settings, showing that CSQ can also evaluate prompt-induced deception.
@@ 6 Conclusion
In this work, we propose a framework to evaluate self-initiated deception in LLMs using two complementary metrics: the Deceptive Behavior Score ($\delta$) and the Deceptive Intention Score ($\rho$). We find that most models exhibit deceptive tendencies that grow with task complexity, and that $\delta$ and $\rho$ are positively correlated, indicating that behavioral inconsistency and strategic intent emerge together. That even advanced models display such systematic deceptive patterns raises serious safety concerns for deploying LLMs in high-stakes decision-making roles.
Reproducibility Statement
We release the code at https://github.com/Xtra-Computing/LLM-Deception. Details of the data and models are provided in Section 5.1 and Appendix A.
@@
Acknowledgement
This research/project is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.
We are grateful to Yue Bi for her helpful suggestions regarding the experimental design in the psychological perspective.
@@ B.2 Deceptive Behavior Analysis
Figures 10 displays the Deceptive Behavior Score ($\delta$) for all evaluated models. The results presented here, which quantify the models’ manifested deceptive actions, are broadly consistent with the analysis in Section 5.2 of the main paper.
@@
Figure 10: Deceptive behavior scores (original, reversed, and geomean) as question scope $n$ varies
@@ B.3 Joint Analysis of Deceptive Intention and Behavior
Since deception is co-determined by deceptive intention and behavior, we study how these metrics change as $n$ increases, with results displayed in Figure 11 supplementary to Figure 5.
@@ C.1 Overall Deceptive Intention
Figure 12: Deception evolution as $n$ increases across all models with 95% confidence intervals. (a) Shows the trend of deceptive intention score, while (b) presents the trend of deceptive behavior score.
@@ C.3 Evolution of Deception by Model Size
In this subsection, we analyze how deception evolves with model size by plotting the deceptive behavior score ($\bar{\delta}$) and the absolute deceptive intention score ($|\bar{\rho}|$) against the number of parameters for open-source LLMs. As shown in Figure 13, both the deceptive behavior score and the deceptive intention score tend to decrease with model size, indicating that scaling generally improves LLM honesty. However, this trend is not particularly strong, as their associated $R^{2}$ values are only around 0.336 and 0.360, respectively. The detailed relationship between model size and LLM deception therefore requires further investigation.
Another observation across different models within the same family is that the effect of model updates on deception is inconsistent: in some series, deception decreases, while in others it increases. For example, within the Qwen family, Qwen3 exhibits a lower trend deceptive behavior score than Qwen2.5, whereas the update from GPT-4o to GPT-4.1 (Figure 6) leads to an increase in both deceptive behavior and deceptive intention scores. These discrepancies likely arise from differences in the LLM training pipelines across companies, which are not fully transparent. As a result, the main factors driving LLM deception remain inconclusive and require further investigation in future model updates.
@@ D.2 Effect of Initial-Followup Difficulty Ratio
As shown in the figure, while the absolute deceptive behavior scores fluctuate with $k$, our key observation is that the relative ranking of the LLMs remains remarkably consistent across the entire range of tested values. This stability is crucial, as it indicates that our evaluation protocol is robust and measures an intrinsic deceptive tendency of the models, rather than an artifact of a specific hyperparameter setting. A consistent ranking validates that our method facilitates a fair comparison among the different models.
Each choice of $k$ corresponds to evaluating deceptive behavior under a different condition. In principle, the optimal approach would be to exhaust all possible $k$ values and aggregate them into an overall $\delta$, but this is computationally prohibitive. Empirically, since the trends are relatively stable across $k$, we instead compare all LLMs under a single $k$ value. The choice of $k$ controls a trade-off between sensitivity to deception and discriminability across models. A larger $k$ induces simpler follow-up questions $n^{\prime}$ that deviate more from the original questions, making any deceptive behavior easier to detect and thus generally increasing $\delta$. However, this also reduces our ability to distinguish models in terms of honesty, as many models reach similarly high $\delta$ in this regime. Conversely, a smaller $k$ focuses the evaluation on a small gap between “belief” and “behavior”; in this case, fewer LLMs exhibit large $\delta$ (e.g., $\delta$ for phi-4 and Llama-3.1 is much smaller), which improves cross-model discriminability. Therefore, to better compare deception across LLMs, we select $k=2$ as a representative setting for all subsequent experiments.
@@
D.3 Variance of Responses to Rephrased Questions
To address potential concerns regarding model sampling randomness, this section details the response variability across rephrased questions. We emphasize that our evaluation’s validity does not require models to have low variance on these questions. Instead, our framework is designed to be robust to such variability. Specifically, all of our main evaluation metrics are presented with 95% confidence intervals (e.g., $\delta$ and $\rho$ in Figure 4 and 5). These intervals are calculated using bootstrapping, a standard statistical method that precisely captures the impact of variance.
Figure 16 presents the detailed Yes/No answer ratios that illustrate this variability. We observe that strong models (e.g., Qwen3, Gemini-2.5-pro, o4-mini) tend to provide highly consistent answers to rephrased questions (nearly all “Yes” or “No”). Conversely, weaker models (e.g., Llama-3.1) provide less certain answers, with ratios closer to 1:1. This observed variability is explicitly and robustly accounted for by the confidence intervals in all other experiments.
Figure 16: Ratio of answers in rephrased questions for each LLM ($Q_{L}$: linked-list question, $Q_{B}$: followup of broken-linked-list question)
@@ F.2 Visualization of Embeddings
This section visualizes how inconsistent responses, which implies deceptive behavior, emerge and cluster within the model’s internal representations as task difficulty ($n$) increases. Figures 22, 23, and 24 display PCA-reduced embeddings, from which we draw two key findings.
@@
Appendix G Generalization to Other Domains
Although CSQ is a simplified, specific question format, LLM’s deception on CSQ strongly suggests the potential for deception in other domains. The structure of CSQ—facts, rules, and question—closely resembles domains such as mathematical proof and logical reasoning. In mathematics, the facts correspond to assumptions, the rules to theorems, and the final question of contactness asks whether a statement can be proved. In other scientific domains, the facts can be experimental results, the rules are physical laws, and the final question is whether a phenomenon can occur. Thus, although CSQ is specific, it resembles a broad range of question types. Extending the evaluation of deception to other domains is therefore a promising direction.
However, a significant challenge is to eliminate the effect of LLM prior knowledge. In the evaluation on practical math questions, the LLM has already learned many implicit rules and facts. Therefore, LLMs may reason according to their internal knowledge rather than relying solely on the provided facts. CSQ mitigates this by using hypothetical names and contactness as facts to avoid triggering internal knowledge. Extending such a framework to other domains such as science, coding, and mathematics is an important direction and will require substantial additional effort.
Appendix H Comparison to Prior Benchmarks and Framework Advantages
Previous benchmarks (Chen et al., 2025; Huang et al., 2024; Wu et al., 2024) have predominantly focused on hallucination and bias, emphasizing performance asymmetry across scenarios that should, in principle, yield similar outcomes. This evaluation on performance asymmetry aligns with our evaluation of deceptive intention in this paper. A more detailed comparison of these works is provided in Table 5. More recently, deception has emerged as another critical concern. Unlike hallucination or bias, deception requires not only asymmetric performance but also self-inconsistency (termed deceptive behavior in this paper). This newly emergent phenomenon has attracted increasing attention and calls for new benchmarks.
While some existing benchmarks evaluate LLM deception, they predominantly focus on prompt-induced behavior, which falls into three main categories. (1) Prompt-Induced Obedience measures compliance with a deceptive persona but conflates obedience with intention (Zhang et al., 2023; Perez & Ribeiro, 2022; Parrish et al., 2021). It only tests if a model can follow a deceptive prompt, not if it would deceive spontaneously. (2) Strategic Game-Playing observes emergent deception in complex, multi-agent games, but these observational findings can suffer from low interpretability (Feng et al., 2024; Chi et al., 2024; Bakhtin et al., 2022). (3) Truthfulness Benchmarks measure a model’s propensity to repeat data-induced falsehoods, which captures epistemic failure as a mistaken belief but not strategic deception (Hong et al., 2024; Lin et al., 2021; Parrish et al., 2021; Malberg et al., 2025). Nonetheless, all these benchmarks evaluate deception by injecting explicit prompts that may impose an implicit objective. Such settings are rarely encountered in everyday use. In contrast, deception under benign prompts—without any explicit contents that could induce an implicit objective—remains largely unexplored.
In contrast, our work shifts from merely observing deceptive outcomes to disentangling deceptive mechanisms using a self-contained logical task with adjustable difficulty (CSQ), yielding three key advantages: Theoretical Grounding. We theoretically separate deceptive intention ($\rho$) from deceptive behavior ($\delta$), providing a direct computational operationalization of the psychological definition of deception (Masip Pallejá et al., 2004), where $\rho$ quantifies a performance asymmetry and $\delta$ quantifies self-consistency. Interpretability. The $\delta$ score localizes failure as belief–expression inconsistency, distinguishing true deception (correct belief, false expression) from hallucination, while $\rho$ serves as an indicator of how a model deceives by revealing its structural preference over deceptive strategies, distinguishing deception from inaccuracy. Robustness. By treating task difficulty as an explicit control knob, our framework stress-tests models under increasing cognitive load and traces $\rho$ and $\delta$ as functions of difficulty, robustly exposing failure points for a variety of LLMs with different scales and capabilities.
Table 5: Comparison of Our Work (CSQ) against prior benchmarks. Performance Asymmetry refers to the divergence in performance between tasks of equivalent difficulty. Self-Consistency refers to the inconsistency between internal belief and external expression. Benign Prompt indicates whether the prompt does not contain any explicit contents that can lead to bias/hallucination/deception.
Benchmark Name & Citation Focus Performance Asymmetry Self- Consistency Benign Prompt
CrowS-Pairs (Nangia et al., 2020) Bias
StereoSet (Nadeem et al., 2020) Bias
BBQ (Parrish et al., 2021) Bias
WinoBias (Zhao et al., 2018) Bias
BOLD (Dhamala et al., 2021) Bias
HolisticBias (Smith et al., 2022) Bias
RealToxicity (Gehman et al., 2020) Bias
MARBLE (Ahmed et al., 2025) Bias
OccuGender (Chen et al., 2025) Bias
TrustLLM (Huang et al., 2024) Bias
TruthfulQA (Lin et al., 2021) Hallucination
HaluEval (Li et al., 2023) Hallucination
FactScore (Min et al., 2023) Hallucination
SAFE (Wei et al., 2024b) Hallucination
SimpleQA (Wei et al., 2024a) Hallucination
FACTOR (Muhlgay et al., 2024) Hallucination
HallusionBench (Guan et al., 2023) Hallucination
FreshQA (Vu et al., 2024) Hallucination
HalluLens (Bang et al., 2025) Hallucination
RAGTruth (Niu et al., 2024) Hallucination
CAP (Gamba et al., 2025) Hallucination
RBench (Wu et al., 2024) Hallucination
Ward et al. (Ward et al., 2023) Deception
Yang et al. (Yang & Buzsaki, 2024) Deception
DarkBench (Kran et al., 2025) Deception
MASK (Ren et al., 2025) Deception
Alignment faking (Greenblatt et al., 2024) Deception
Sandbagging (Van Der Weij et al., 2024) Deception
DeceptionBench (Ji et al., 2025) Deception
Diplomacy scenario (Park et al., 2024) Deception
Semantic triggers (Hagendorff, 2024) Deception
AmongUs (Chi et al., 2024) Deception
Human-level games (Bakhtin et al., 2022) Deception
Our Work (CSQ) Deception
Appendix I Broader Impact